you can see a local tv on summer morning temparature will rise from the current 18degree celsius-39 degree celsius.the corresponding increase in the foren heat temperature=

Grade:12

1 Answers

Bejoi Mathew
7 years ago
Farenheit temperature to Celcius Temperature conversion technique is as follows:
F= 9/5 C + 32
Plug in the values you are sure to get the answer.
example,
F1= 9/5 * 18 + 32
F2= 9/5 * 39 + 32
F1 to F2 is the corresponding increase in farenheit scale.
HOPE THIS HELPS
THANK YOU